A model aircraft is flying north with a velocity of 24m/s. A wind is blowing from west to east at 10m/s. What is the resultant velocity of the plane ?
1
Expert's answer
2018-09-17T15:25:08-0400
The magnitude of resultant velocity obtained from the Pythagorean theorem
v^2=〖24〗^2+〖10〗^2
v=√(〖24〗^2+〖10〗^2 )=√(2^2⋅〖12〗^2+2^2⋅5^2 )=2⋅√(〖12〗^2+5^2 )=2⋅√(144+25)=2⋅√169=2⋅13=26 m⁄s